# White Men Lead, Black Women Help: Uncovering Gender, Racial, and Intersectional Bias in Language Agency

### v1

@@ Abstract
Social biases can manifest in language agency. For instance, White individuals and men are often described as “agentic” and achievement-oriented, whereas Black individuals and women are frequently described as “communal” and as assisting roles. This study establishes agency as an important aspect of studying social biases in both human-written and Large Language Model (LLM)-generated texts. To accurately measure “language agency” at sentence level, we propose a Language Agency Classification dataset to train reliable agency classifiers. We then use an agency classifier to reveal notable language agency biases in $6$ datasets of human- or LLM-written texts, including biographies, professor reviews, and reference letters. While most prior NLP research on agency biases focused on single dimensions, we comprehensively explore language agency biases in gender, race, and intersectional identities. We observe that (1) language agency biases in human-written texts align with real-world social observations; (2) LLM-generated texts demonstrate remarkably higher levels of language agency bias than human-written texts; and (3) critical biases in language agency target people of minority groups—for instance, languages used to describe Black females exhibit the lowest level of agency across datasets. Our findings reveal intricate social biases in human- and LLM-written texts through the lens of language agency, warning against using LLM generations in social contexts without scrutiny.
White Men Lead, Black Women Help: Uncovering Gender, Racial, and Intersectional Bias in Language Agency
Yixin Wan and Kai-Wei Chang {elaine1wan, kwchang}@cs.ucla.edu
@@ 1 Introduction
Social biases manifest through different levels of agency in texts describing individuals of different demographic groups, such as race and gender Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023). This is often through portraying minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020).
@@
To measure agency level in texts, previous works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022) mostly utilized string matching for words in agentic and communal lexicons. However, existing tools like string matching and sentiment classifiers often fail to capture language agency. Table 11 in Appendix B shows that string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively. Table 1 provides a qualitative example: while differences in language agency are clearly observable in the texts, string matching yields 00 agentic and 00 communal words for both texts; sentiment classifier labels both sentences as “positive”, failing to identify implicit bias in agency.  Wan et al. (2023) utilized a model-based method for measuring agency and related gender biases, but only in the context of reference letters. Most aforementioned studies also only focused on single dimensions of biases and single types of text, limiting the scope of analysis.
Occupation Biography String Match Sentiment Classifier Agency Classifier
# Agen. # Comm.
Male Software Engineer Growing up in Vienna, he previously co-founded a legal marketplace startup connecting attorneys with potential clients, and has accumulated over 10 years of experience in iOS security research and app development, leading development on several high-profile projects such as … 0 0 Positive Agentic
Female Software Engineer She is passionate about helping developers build engaging, useful experiences and services that make people‚ lives better. 0 0 Positive Communal
Table 1: Example of gender biases manifested in language agency. Agentic languages are highlighted in blue, and communal languages are highlighted in red. String matching method (columns 3 and 4) and sentiment-based approach (column 5) fail to capture the bias in language style, whereas our proposed language agency classifier successfully identifies gender bias in agency.
Our study refines and extends the model-based approach to capture language agency biases in a broader scope of human-written and LLM-generated texts. In addition to studying both gender and racial biases, we propose to investigate intersectional biases in language agency. We first contribute a sentence-level Language Agency Classification dataset for training effective agency classifiers. Utilizing an agency classifier trained on our dataset, we reveal remarkable gender, racial, and intersectional biases in language agency on $6$ human-written and LLM-synthesized datasets, including biographies, professor reviews, and reference letters.
Results show that: (1) our findings on biases in human-written texts agree with previous social science discoveries. For instance, biographies of individuals with male-dominant occupations—such as pastor, architect, and software engineer—demonstrate greater extents of gender bias. This proves language agency as a reliable tool for studying biases in texts. (2) LLM-generated texts demonstrate notably greater levels of language agency gender bias, compared to human-written texts. (3) Both human-written and LLM-generated texts demonstrate remarkably higher language agency when describing dominant social groups—such as White men—compared to texts about minority groups—such as Black women, revealing severe biases in language agency. For example, reviews for Black professors—especially Black female professors—have the lowest language agency levels among faculties of all races. Black individuals also receive recommendation letters with the lowest language agency level, while White people—especially White males—receive the most agentic letters of recommendation. This study unveils critical language agency biases in human- and LLM-written biographies, professor reviews, and reference letters. Our findings warn against using model-generated texts while overlooking potential risks and harms in society.  [^1] Dataset and model checkpoints will be released upon acceptance.
@@ 2 Related Work
2.1 Biases in Human-Written Texts
The widespread presence of gender, racial, and intersectional bias in human society has significantly impacted human language Blodgett et al. (2020); Doughman et al. (2021). Since generative LLMs use extensive texts for training, understanding biases in texts is crucial for identifying and preventing biases in language technologies.
Bias in BiographiesWagner et al. (2016); Field et al. (2022), and  Park et al. (2021) studied gender biases in Wikipedia biographies.  Park et al. (2021) analyzed biases in power, agency, and sentiment words in biography pages;  Wagner et al. (2016) revealed negative linguistic biases in womens’ pages.  Field et al. (2022) and  Adams et al. (2019) studied racial biases in editorial traits such as length and academic rank.  Field et al. (2022); Adams et al. (2019) and  Lemieux et al. (2023) stressed the importance of studying intersectional gender and racial biases in Wikipedia. Along similar lines, Otterbacher (2015) found that IMDB biographies for Black female actresses describe them with more heavily subjective adjectives.
Bias in Professor ReviewsSchmidt visualized the extent of gendered language in RateMyProfessor reviews by string matching for gender-indicative words. Their observations align with prior findings on gender biases in student ratings for professors, revealing that instructors with female perceived gender received lower ratings than males (Roper, 2019; Macnell et al., 2014).  Reid (2010) showed that professors from racial minority groups—specifically Black and Asian groups—received more negative RateMyProfessor evaluations on overall quality than White faculties.  Chávez and Mitchell (2020) explored intersectional gender and racial biases in professor reviews, and found that female professors from racial minority groups received lower student evaluation scores compared to white males.
@@ 2.2 Bias in Language Agency
An increasing body of recent studies have investigated biases in intricate language styles, such as language agency Sap et al. (2017); Ma et al. (2020); Stahl et al. (2022); Wan et al. (2023).
Agentic language describes “proactive” characteristics such as speaking assertively, influencing others, and initiating tasks. Communal language portrays “reactive” characteristics like caring for others, providing assistance, and sustaining relationships (Madera et al., 2009; Wan et al., 2023). Bias in language agency can arise due to the disproportionate portrayal of agency and power, e.g. by showing women as submissive and powerless and men as assertive and dominant (Stahl et al., 2022), or by describing racial minority groups with more communal language than agentic (Grimm et al., 2020; Polanco-Santana et al., 2021).
@@
Evaluation    Akos and Kretchmar (2016); Sap et al. (2017); Ma et al. (2020); Grimm et al. (2020); Polanco-Santana et al. (2021); Park et al. (2021), and Stahl et al. (2022) measured language agency by string matching for agentic and communal verbs, and then calculate their occurrence frequencies. However, string-matching methods fail to consider the diversity and complexity of language, and could not capture implicit indicators of language agency, as illustrated in Figure 1.  Wan et al. (2023) was the first to adopt a model-based method to measure language agency gender biases in LLM-generated reference letters. Nevertheless, the scope of their analysis is constrained to the LLM-synthesized letters of recommendation.
We believe that language agency is a common theme over a broad scope of social biases across all types of human-written and model-generated texts. Therefore, we propose to explore gender, racial, and intersectional biases in language agency in a broader context.
3 Language Agency Bias
3.1 Definition
We define “language agency bias” to be the unequal representation of language agency in texts depicting people of different demographic identities, such as race and gender. Given a piece of natural language text, we calculate the agency of each sentence, and then measure language agency level as the percentage of agentic sentences within the text. Biases can then be manifested as harmful differences between the agency levels of texts about people of different genders or racial groups. Different from previous works (Sap et al., 2017; Ma et al., 2020; Stahl et al., 2022), which used string matching for agentic and communal words from pre-defined lexicons, we adopt a model-based approach through a language agency classification model, which can more accurately capture implicit language agency traits in texts. Different from Wan et al. (2023), we propose a sentence-level language agency classification dataset, and extend the scope of analysis for language agency bias to a broader scope of natural language texts, including personal biographies, professor reviews, and reference letters.
3.2 Language Agency Classification Dataset
We propose the Language Agency Classification (LAC) dataset, a ChatGPT (OpenAI, 2022)-generated and human-reannotated dataset, to build reliable and accurate agency classifiers for identifying biases in language agency.
@@ 3.2.1 Dataset Collection
To collect agentic and communal sentences, we first preprocess a personal biography dataset (Lebret et al., 2016) into sentences, then use ChatGPT to paraphrase each sentence into an agentic version and a communal version. Next, we invite $2$ human annotators to re-label each data as “agentic”, “communal”, or “neutral”. Full details for constructing the LAC dataset are in Appendix A.
@@ 3.2.2 Dataset Statistics
Inter-Annotator Agreement   We consider ChatGPT’s categorization of agentic and communal paraphrases as its classification labels. Then, we calculate Fleiss’s Kappa score Feinstein and Cicchetti (1990) between ChatGPT’s labeling and the two main human annotators. The reported Fleiss’s Kappa score is $0.71$, proving the satisfactory quality of our proposed LAC dataset. Observing a significantly smaller number of “neutral” entries compared to the $2$ other labels, we remove all “neutral” entries in the dataset, achieving a new Fleiss’s Kappa score of $0.90$ for the final LAC dataset.
@@
Dataset Split   We split the annotated and aggregated dataset into Train, Test, and Validation sets with a $0.8$, $0.1$, $0.1$ ratio. Detailed statistics of each split are in Appendix A, Table 9.
4 Experiments
4.1 Training Language Agency Classifiers
@@
We experiment with both discriminative and generative models as base models for training language agency classifiers. For discriminative model, we experiment with BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019). For generative model, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. (2023). Appendix B provides details of training and inferencing the classifiers, in which Table 11 reports classifier performances. Based on model performance on LAC dataset’s test set, we choose the fine-tuned BERT model as the language agency classifier in further experiments.
4.2 Revealing Language Agency Biases
4.2.1 Datasets
We experimented with $6$ datasets—including biographies, professor reviews, and reference letters—to reveal gender, racial, and intersectional biases through the lens of language agency. $3$ of the datasets were publicly accessible; due to a lack of data with racial information, we use ChatGPT to synthetically generate the rest $3$ of the datasets with racial information. Details of all datasets are in Appendix C, Table 12.

### v3

# White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs

@@ Abstract
Language agency is an important aspect of evaluating social biases in texts. While several studies approached agency-related bias in human-written language, very limited research has investigated such biases in Large Language Model (LLM)-generated content. In addition, previous research often relies on string-matching techniques to identify agentic and communal words within texts, which fall short of accurately classifying language agency. We introduce the novel Language Agency Bias Evaluation (LABE) benchmark, which comprehensively evaluates biases in LLMs by analyzing agency levels attributed to different demographic groups in model generations. LABE leverages 5,400 template-based prompts, an accurate agency classifier, and corresponding bias metrics to test for gender, racial, and intersectional language agency biases in LLMs on 3 text generation tasks: biographies, professor reviews, and reference letters. To build better and more accurate automated agency classifiers, we also contribute and release the Language Agency Classification (LAC) dataset, consisting of 3,724 agentic and communal sentences. Using LABE, we unveil previously under-explored language agency social biases in 3 recent LLMs: ChatGPT, Llama3, and Mistral. We observe that: (1) For the same text category, LLM generations tend to demonstrate higher levels of gender bias than human-written texts; (2) On most generation tasks, models demonstrate remarkably higher levels of intersectional bias than the other bias aspects. Those who are at the intersection of gender and racial minority groups—such as Black females—are consistently described by texts with lower levels of agency, aligning with real-world social inequalities; (3) Among the 3 LLMs investigated, Llama3 demonstrates greatest overall bias in language agency; (4) Not only does prompt-based mitigation fail to resolve language agency bias in LLMs, but it frequently leads to the exacerbation of biases in generated texts.
@@ 1 Introduction
Social biases manifest through the varying levels of agency in texts describing different demographic groups Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023). For instance, bias exists in texts portraying demographic minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020). While a body of works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022; Wan et al., 2023) have studied agency level in texts, these previous works suffer from several remarkable drawbacks:
1. Existing works fail to establish a comprehensive evaluation benchmark for language agency biases in LLMs. Most aforementioned studies studied such biases in single types of human-written texts, and only focused on single dimensions of bias (e.g. only gender bias), limiting the scope of their analysis. As people are exploring more real-world downstream applications of LLM-generated texts, it is critical to identify and quantify potential agency-related fairness issues in LLM generations.
@@
2. Existing methods to measure language agency struggle with achieving accuracy and reliability. Prior works often utilized string matching for words in agentic and communal lexicons to measure agency. However, string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively (as shown in Appendix C, Table 11). A qualitative example is provided in Figure 1: while differences in language agency are clearly observable in the texts, string matching yields 00 agentic and 00 communal words for both texts; sentiment classifier labels both sentences as “positive”.  Wan et al. (2023) utilized a model-based method for measuring agency, but their model only achieves $66.49\%$ classification accuracy (Appendix C, Table 11).
Figure 1: Example of using LABE to measure bias in biography generation. Agentic and communal phrases are highlighted in blue and red. Despite the obvious bias, prior methods (string matching, sentiment-based) fail to capture differences. LABE’s agency classifier successfully identifies the bias.
To address these research gaps, our study proposes a novel Language Agency Bias Evaluation (LABE) benchmark for comprehensively measuring gender, racial, and intersectional language agency biases in LLMs. Using 5,400 template-based entries, an accurate language agency classifier, and interpretable metrics for each bias dimension, LABE examines agency-related biases on 3 text generation tasks for LLMs: biography, professor review, and reference letter generation. For building the accurate and reliable automated agency classification tool, we also collect and contribute the Language Agency Classification (LAC) dataset. Using a scalable, consistent, natural, and fair data collection pipeline with LLM-empowered data generation and human cross-verification, we constructed LAC with 3,724 agentic and communal sentences. Finally, we trained an agency classifier with LAC (achieving 91.69% test accuracy) and incorporated it into LABE to evaluate language agency biases in 3 recent LLMs: ChatGPT, Mistral, and Llama3. We observed that:
  • LLMs show greater language agency bias than humans. For the same text type (e.g. reference letter), LLM generations are often more gender-biased than human-written texts.
  • Language agency biases target intersectional minority groups. For instance, Black professors—especially Black female professors—have the lowest language agency levels among faculties of all races in ChatGPT and Llama3-generated professor reviews.
  • Llama3 appears to be most biased among the 3 LLMs investigated. Llama3 possesses the greatest overall level of language agency bias; most severe biases in the model are observed in professor review and reference letter generation.
  • Simple prompt-based mitigation methods might exacerbate biases. Contrary to expectations, instructing the model on avoiding biases fails to resolve the fairness issue. Moreover, it oftentimes results in even higher levels of bias in LLM-generated texts.
Our LABE benchmark and LAC dataset introduce novel and valuable technical contributions, and our findings reveal previously unexplored fairness risks of LLMs from the perspective of language agency. Furthermore, we unveil the shocking fact that widely adopted prompt-based mitigation methods may intensify language agency biases in LLMs; more effective methodologies need to be developed to address the complex fairness challenge.
@@ 2 Related Work
2.1 Biases in Human-Written and LLM-generated Texts
The presence of gender, racial, and intersectional bias in human society has significantly impacted human language Blodgett et al. (2020); Doughman et al. (2021) and generative LLMs, which utilize extensive texts for training. In this work, we investigate biases in 3 different categories of texts: biographies, professor reviews, and reference letters.
Bias in BiographiesWagner et al. (2016); 10.1145/3485447.3512134, and  Park et al. (2021) studied gender biases in Wikipedia biographies.  Park et al. (2021) analyzed biases in power, agency, and sentiment words in biography pages;  Wagner et al. (2016) revealed negative linguistic biases in womens’ pages.  10.1145/3485447.3512134 and  Adams et al. (2019) studied racial biases in editorial traits such as length and academic rank.  10.1145/3485447.3512134; Adams et al. (2019) and  Lemieux et al. (2023) stressed the importance of studying intersectional gender and racial biases in Wikipedia. Along similar lines, Otterbacher (2015) found biases towards Black female actresses in IMDB biographies. Bias in Professor Reviews   Prior works (Roper, 2019; Macnell et al., 2014) have revealed gender biases in student ratings for professors—instructors with female perceived gender received lower ratings than males. Schmidt visualized the gendered language in RateMyProfessor reviews by string matching for gender-indicative words.  Reid (2010) showed that professors from racial minority groups received more negative RateMyProfessor evaluations.  Chávez and Mitchell (2020) further revealed intersectional gender and racial biases towards female professors from racial minority groups in professor reviews.
@@ 2.2 Bias in Language Agency
An increasing body of recent studies have investigated biases in intricate language styles, such as language agency Sap et al. (2017); Ma et al. (2020); Stahl et al. (2022); Wan et al. (2023).  Akos and Kretchmar (2016); Sap et al. (2017); Ma et al. (2020); Grimm et al. (2020); Polanco-Santana et al. (2021); Park et al. (2021), and Stahl et al. (2022) measured language agency by string matching for agentic and communal verbs, and then calculate their occurrence frequencies. However, string-matching methods fail to consider the diversity and complexity of language, and could not capture implicit indicators of language agency, as illustrated in Figure 1.  Wan et al. (2023) was the first to adopt a model-based method to measure language agency gender biases in LLM-generated reference letters. Nevertheless, their model lacks accuracy in sentence-level classification, and the scope of their analysis is constrained to LLM-synthesized reference letters.
3 The Language Agency Bias Evaluation (LABE) Framework
Agentic language depicts “proactive” characteristics such as speaking assertively, influencing others, and initiating tasks; communal language portrays “reactive” characteristics like caring for others, providing assistance, and sustaining relationships (Madera et al., 2009; Wan et al., 2023). We define “language agency bias” to be the unequal representation of language agency in texts depicting different demographic groups, e.g. by showing women as submissive and powerless and men as assertive and dominant (Stahl et al., 2022), or by describing racial minority groups with more communal language than agentic (Grimm et al., 2020; Polanco-Santana et al., 2021).
In this paper, we propose the Language Agency Bias Evaluation (LABE) benchmark for comprehensively assessing language agency biases in LLMs across race, gender, and intersectional identities. LABE prompts an LLM to generate descriptive texts for multiple demographic groups, and assesses biases by inspecting the variability in language agency between generations for different groups.
Split # Entries # Agentic # Communal
Train $2,979$ $1,523$ $1,456$
Valid $372$ $192$ $180$
Test $373$ $199$ $174$
Table 1: LAC Dataset Statistics.
Text Type Template # Entries
Biography “Generate a personal biography for {name}, a {age} year old {race} {gender} {occupation}.” 2,400
Professor Review “Generate a review for {name}, a {race} {gender} professor in the {department name}.” 600
Reference Letter “Generate a detailed reference letter for {name}, a {age} year old {race} {gender} {occupation}.” 2, 400
Table 2: Templates for LLM Generations in LABE.
3.1 Generative Discriptive Texts for Demographic Groups with LLMs
Wan et al. (2023) proposed the Context-Less Generation (CLG) setting, in which they adopt templates and descriptors to prompt for a variety of LLM-generated reference letters for different genders. Inspired by CLG, we extend the setting to 3 different text generation tasks: biography, professor review, and reference letter generation. We combine descriptors with demographic information—such as race, gender, or intersectional identities—and template-based prompts to query for LLMs’ generation. Each prompt must contain race and gender descriptors. For the name descriptor, we prompt ChatGPT to generate 5 popular names for each gender and race intersectional group. Descriptors for additional details like occupation and department are included to improve prompt variability. The final LABE benchmark tests LLMs on 2,400 templated-based prompts for biography generation, 600 for professor review, and 2,400 for reference letters. Note that entry numbers differ due to the difference in descriptors used (departments for professor review, whereas occupations for the other 2). We provide examples of each descriptor below, with full details included in Appendix A:
  1. 1.
    Race: “Black”, “White”, “Hispanic”, “Asian”
  2. 2.
    Gender: “Male”, “Female”
  3. 3.
    Name:
    1. (a)
      “Asian” + “Female”: “Mei”, “Aiko”, “Linh”, “Priya”, “Ji-Yoon”
    2. (b)
  4. 4.
    Occupation: “student”, “entrepreneur”,“actor”, “artist”, “chef”, …
  5. 5.
    Department: “Communication department”, “Fine Arts department”, “Chemistry department”, …
3.2 Evaluating Language Agency: The Language Agency Classification (LAC) Dataset
For building accurate automated evaluation tools for language agency, we propose the Language Agency Classification (LAC) dataset, a corpus with 3,724 agentic and communal sentences with corresponding labels. The dataset construction process incorporates an efficient automated generation pipeline, and careful verification by human annotators who are native speakers of English.
@@ 3.2.1 Dataset Collection
To ensure the trustworthiness of the constructed dataset, we identified and followed 4 core pillars for the data collection process:
  1. 1.
    The data construction process should be scalable. Since we are constructing a classification dataset, we need to ensure enough entries for agentic language and communal language to train a useful classifier.
  2. 2.
    The data construction process should be consistent. Quality of texts for all entries should remain consistent, and should not differ from part to part.
  3. 3.
    The data construction process should be natural. We need to ensure that data labels align with human perceptions of ’agentic’ and ’communal’ language.
  4. 4.
    The data construction process should be fair. Balancing measures should be taken to prevent potential biases from label imbalance. If the constructed data is built based on an original dataset, we also need to ensure there is no social bias propagation from the original data.
To collect a dataset with agentic and communal sentences through a mechanism that is scalable, consistent, natural, and fair, we adopt a novel dataset construction framework that consists of an automated component and a human-involved component. We begin by preprocessing a personal biography dataset (Lebret et al., 2016) into sentences, aiming at using these as seed texts to construct agentic and communal texts through paraphrasing. This step ensures the fairness of collected dataset, since (1) the raw data output would be balanced between the two labels, and (2) each sentence in each biography would have an agentic paraphrase and a communal paraphrase, preventing social bias propagation like having more agentic sentences for dominant social groups. Next, we adopt Openai’s gpt-3.5-turbo-1106 model (OpenAI, 2022) to paraphrase each sentence into an agentic version and a communal version. This ensures scalability through an automated generation pipeline, and also guarantees consistency since all paraphrases would come from a single source (in contrast with using human-written paraphrases, which is hard to scale and might result in drastically subjective writing tones). Furthermore, we utilize a human verification step to ensure the naturalness of the generated dataset. We invite $2$ human annotators, who are native speakers of English, to re-label each data and identify ambiguous cases. Finally, data entries with ambiguity are removed and ground truth labels of the LAC dataset are decided by a majority vote between the annotators’ labels and the paraphrasing target (i.e. whether a sentence was generated as an “agentic” or “communal” paraphrase). We elaborate on full details of dataset construction in Appendix B.
@@ 3.2.2 Dataset Statistics
The finalized LAC dataset consists of 3,724 entries. Below, we present the data statistics.
@@
Inter-Annotator Agreement   We consider the paraphrasing target—whether a text was generated to be “agentic” or “communal”—as the default labels from the automated paraphrasing pipeline. Then, we calculate Fleiss’s Kappa score Feinstein and Cicchetti (1990) between the default labels and the two main human annotators. The finalized version of the proposed LAC dataset achieves a Fleiss’s Kappa score of 0.90, proving the satisfactory quality of the dataset.
@@
Dataset Split   To adapt the constructed dataset for training and inferencing language agency classifiers, we split the annotated and aggregated dataset into Train, Test, and Validation sets with a $0.8$, $0.1$, $0.1$ ratio. Detailed statistics of each split are in Table 2.
3.2.3 Building A Language Agency Classifier With LAC
@@
We experiment with both discriminative and generative models as base models for training language agency classifiers. Based on performances on LAC’s test set, we choose the fine-tuned BERT model as the language agency classifier in further experiments. Appendix C provides details of training and inferencing the classifiers, in which Table 11 reports classifier performances.
3.3 Quantifying Language Agency Bias in LLMs
We use the LAC-trained agency classifier to build quantitative metrics for measuring language agency bias in LLM generations. Specifically, we designed $2$ types of metrics: Intra-Group Agentic-Communal Ratio Gaps and Inter-Group Ratio Gap Variances. Intra-Group metrics objectively measure the agency level in texts generated for different demographic groups, whereas inter-group metrics estimate the variability of agency levels across groups.
Intra-Group: Ratio Gaps between Agentic and Communal Sentences. For a piece of LLM-generated text, we first calculate the average percentage of agentic and communal sentences. We then report the intra-social-group average ratio gap between agentic and communal sentences to better reflect the absolute level of language agency.
Inter-Group: Variance of Ratio Gaps.   We also design inter-group metric that reflect biases through relative agentic level differences between social groups. To better estimate the variability of bias levels across multiple groups (e.g. intersectional gender and racial identities), we mainly report the variance of the agentic-communal ratio gaps across all demographic groups.
4 Unveiling Language Agency Biases in LLMs with LABE
We utilize LABE to conduct experiments on measuring gender, racial, and intersectional biases in 3 recent LLMs: ChatGPT (OpenAI, 2022), Llama3 (Touvron et al., 2023), and Mistral (AI, 2023).
Models and Generation Settings   We experiment with 3 recent LLMs: the gpt-3.5-turbo-1106 version of OpenAI s ChatGPT (OpenAI, 2022), Llama3-8B-Instruct (Touvron et al., 2023), and Mistral-7B-Instruct-v0.2. We utilize ChatGPT’s API for experiments, with no license information. Llama3 is licensed under the Meta Llama 3 Community License and Mistral is under Apache License 2.0; both models are publicly available. For ChatGPT, we followed all default generation settings in the API call. We use Huggingface’s text generation pipeline to implement Llama3 and Mistral, and follow all default generation hyperparameters besides setting maximum number of new tokens to 512. All results are averaged on random seeds 00, $1$, and $2$.
4.1 Findings 1: LLM generations are More Gender Biased than Human-Written Texts
We establish comparison with bias in LLM-generated texts by incorporating analysis on 3 existing datasets: human-written biographies in Bias in Bios, human-written professor reviews on RateMyProfessor, and the reference letter dataset in  Wan et al. (2023)’s work, which consists of letters generated by LLMs given extensive biographical information (e.g. multi-sentence descriptions of career development) about specific individuals. Since we do not find any publicly available large-scale dataset for reference letters,  Wan et al. (2023)’s data is our best choice as a proxy of human-written letters. In addition, we did not find any openly-accessible datasets from the 3 categories that include racial information, limiting our analysis to gender biases.
Figure 2: Visualization of language agency gender bias in human-written and LLM-generated texts. Y-axis denotes the gender differences in agentic-communal ratio gaps in texts (ratio gap in male texts - female texts). On all types of texts, LLM demonstrates greater bias than humans.
Dataset Model Gender Diff. (M-F)
Biography Human $10.12$
ChatGPT $8.49$
Mistral 10.87
Llama3 $8.51$
Professor Review Human $1.86$
ChatGPT $6.57$
Mistral $8.14$
Llama3 $11.51$
Reference Letter Wan et al. (2023) $4.64$
ChatGPT $9.33$
Mistral $10.84$
Llama3 $9.44$
[…]
table Language agency gender bias in human-written and LLM-generated texts, measured by gender difference in agency-communal ratio gaps. Highest bias for each type of text is in bold.
4.1.1 Human-Written Texts: Dataset Details
We experiment with $3$ publicly accessible datasets of personal biographies, professor reviews, and reference letters. Full details of all datasets are in Appendix D.